# Quantum decoherence by magnetic fluctuations in a candidate axion insulator

### v1

@@ Abstract
A hallmark of second-order phase transitions is the phenomenon of “critical slowing down,” a significant increase in the time it takes for a system to respond to perturbations or to return to equilibrium. In the critical regime, the free energy is flat and the dynamics of the system become extremely slow. Within the context of a magnetic transition, this occurs when locally-fluctuating moments become entangled with their neighbors at increasingly large distances. Both the correlation length $\xi\propto\left|T-T_{c}\right|^{-\nu}$ and the correlation time $\tau\propto\xi^{z}$ diverge, with critical exponents $\nu$ and $z$, respectively [33]. In an ideal system, this results in $\gamma\rightarrow 0$ at the critical temperature, which we propose as the likely origin for the sharp peak in SHG that we observe at $T_{N}$. We emphasize that within this interpretation, the growth of SHG as temperature approaches $T_{N}$ is the result of an increase in the lifetime of magnetic fluctuations, not an increase in their amplitude. We can place a rough upper bound on the coherence time $\tau$ of the $4f^{7}\rightarrow 4f^{6}$ magnetic excitation caused by the charge-transfer optical transition. We observe an $O(1)$ change in SHG over the measured temperature range, which implies that $\gamma$ must be at least of the same order as $\Delta\omega$, the spectral width of the ultrafast laser. Thus, we calculate $\tau\sim 1/\gamma<1/\Delta\omega\sim 16$ fs. This value is comparable to the excitation lifetime of an isolated magnetic atom bound to a metallic substrate, as measured by inelastic tunneling spectroscopy, where it was suggested that the short lifetime was due to hybridization with metallic bands, leading to efficient electron-electron interactions that relax the magnetic state of the atom [34]. EuSn2(As,P)2 is also metallic, and indeed it has been suggested that the magnetic exchange coupling in the material is mediated by conduction electrons via the Ruderman–Kittel–Kasuya–Yosida (RKKY) interaction [20, 35]. Thus, we may also expect a similarly short magnetic excitation lifetime due to coupling to conduction electrons.
@@
According to the celebrated fluctuation-dissipation theorem, there is a close relationship between the imaginary part of a linear susceptibility and the power spectral density of fluctuations of that susceptibility’s conjugate observable [33]. No such theorem exists for nonlinear susceptibilities, where the imaginary part does not represent dissipation because the time-average of $\mathbf{E}(\omega)\cdot\mathbf{J}(2\omega)$ is necessarily zero [36]. It has been theoretically shown, however, that nonlinear responses are much more sensitive than linear responses to semiclassical chaotic behavior through the interference of multiple dynamical Liouville-space trajectories [37]. The second-order susceptibility, for example, is directly related to the stability matrix of a system and its associated Lyapunov exponents quantifying the system’s sensitivity to initial conditions. Our measurements largely affirm this notion, demonstrating that, at least near resonances, the nonlinear optical response of a material can be significantly influenced by the dynamics of fluctuations of some order parameter of the material, even when that order is not the conjugate observable of the susceptibility. We expect that this principle is generally applicable, and that under sufficiently resonant conditions, the nonlinear optical response of a material can be harnessed to study fluctuations, dynamics, and quantum coherence across many types of second-order phase transitions.
@@
Figure 3: Spontaneous rotational symmetry breaking in EuSn2As2. (a) Temperature dependence of the 2-fold rotational anisotropy that emerges at the antiferromagnetic phase transition. The inset shows Sin-Uout (SU) RA-SHG patterns for each temperature, and the main panel shows the corresponding 2-fold anisotropy versus temperature. The gray curve is a guide to the eye, showing an order-parameter-like onset below $T_{N}$. (b) Spatial map of rotational symmetry breaking. RA-SHG patterns where acquired on a $10\times 10$ grid spanning a $300\times 300$ $\mu$m area. At each grid point, the amplitude and direction of the 2-fold anisotropy was extracted (represented by the length and direction, respectively, of the double-headed arrows). The upper left inset shows a photograph of the sample with the mapped region identified, and the lower left inset shows a polar plot of the anisotropy sorted by direction into three symmetry-equivalent antiferromagnetic domains (red, green, and blue). The map reveals static domains with sizes on the order of $\sim$100 $\mu$m.
@@
Within the long-range ordered antiferromagnetic phase below $T_{N}$, the Eu2+ moments of EuSn2(As,P)2 become in-plane aligned. This necessarily results in spontaneous breaking of the $\bar{3}$ symmetry axis of the crystal, as the moments must choose one of three equivalent crystallographic directions along which to orient themselves. Rotational symmetry breaking can generally be revealed through RA-SHG patterns, which are exceptionally sensitive to point group symmetries [26]. As Fig. 3(a) demonstrates, we indeed observe the onset of rotational symmetry breaking below $T_{N}$ in RA-SHG patterns, with a clear order-parameter-like temperature dependence. Here we measure the unpolarized SHG output (Uout) because, in this case, a fully isotropic (i.e. circular) response is expected in the high-symmetry phase, degenerating into a 2-fold “peanut-like” shape when the rotational symmetry axis is lost. Further details on the quantification of the 2-fold anisotropy can be found in the Supplemental Materials. We attribute this observation to two possible mechanisms, both involving spin-orbit coupling. In the first mechanism, strong coupling of the globally-aligned spins to the initial-state $f$ orbitals ($f_{xyz}$, $f_{xz^{2}}$, and $f_{x(x^{2}-3y^{2})}$) shifts their energies, and this manifests as slight changes to the optical resonance condition for different polarization angles. In the second mechanism, the global alignment of moments influences the quantum coherence of different initial-state $f$ orbitals. Given that the 2-fold anisotropy we observe is almost of the same magnitude as the SHG intensity itself ($\sim$20%), we believe the second scenario is more likely. Accompanying spontaneous symmetry breaking is the formation of magnetic domains transforming into each other by the broken symmetry element. Optical imaging of antiferromagnetic domains is typically very challenging because of the absence of a net magnetization, with SHG being one of the few known techniques for doing so [38, 39]. Figure 3(b) shows a spatial RA-SHG mapping over a $300\times 300$ $\mu$m area that reveals these domains. We find a roughly equal distribution of all three possible domain orientations, with an approximate domain length scale of $\sim$100 $\mu$m. Thermal cycling across $T_{N}$ generates completely new domain patterns, as shown in the Supplemental Materials, suggesting domain boundaries are not pinned by disorder.
@@
In conclusion, we have harnessed near-resonant EQ-SHG to probe the intricate magnetic dynamics in the candidate axion insulator EuSn2(As,P)2. By examining the material across its antiferromagnetic phase boundary, we have revealed a pronounced dimensional crossover in the quantum decoherence induced by magnetic fluctuations, with two-dimensional in-plane ferromagnetic correlations at high temperatures transitioning to three-dimensional long-range order below $T_{N}$. Moreover, we have observed an unusually sharp SHG peak right at the magnetic phase boundary that we attribute to the phenomenon of critical slowing down, where the dynamics of the system become extremely slow and the correlation length and time diverge. The quantum coherence time of the charge-transfer $4f^{7}\rightarrow 4f^{6}$ magnetic excitation is estimated to be no more than $\sim 16$ fs. We conjecture that such a short lifetime is due to efficient coupling between the localized magnetic moments and the conduction electrons in EuSn2(As,P)2, supporting a scenario whereby RKKY interactions mediate the magnetic exchange [20, 35]. This conclusion has important implications for the coupling strength between the magnetic order and the topological electronic states, where prior work has reported only a very weak hybridization [22]. Exotic topological boundary modes have been predicted at magnetic domain boundaries in this class of materials [5]. By leveraging the extreme sensitivity of RA-SHG, we have uncovered the clear breaking of rotational symmetry within the antiferromagnetic phase and used it to map out the spatial domain structure that emerges. Our findings help to clarify the nature of the magnetic order in this class of materials and its interactions with topological surface states, which play a critical role in realizing topological magnetoelectric phenomena. More generally, our work showcases the unique capabilities of nonlinear optical spectroscopy as a powerful tool for investigating quantum coherence and fluctuations in magnetic quantum materials.
Acknowledgments

### v2

@@ II Results
A hallmark of second-order phase transitions is “critical slowing down,” a significant increase in the time it takes for a system to respond to perturbations and return to equilibrium. Within the context of magnetic transitions, this occurs when locally-fluctuating moments become entangled with their neighbors at increasingly long distances. Both the correlation length $\xi\propto\left|T-T_{c}\right|^{-\nu}$ and time $\tau\propto\xi^{z}$ diverge at $T_{N}$, with critical exponents $\nu$ and $z$, respectively [33]. In Supplementary Note 4, we derive the relationship between the off-diagonal quantum decoherence rate $\gamma_{nm}$ and the magnetic correlation time: $\gamma^{-1}_{nm}=2\tau$. This results in $\gamma_{nm}\rightarrow 0$ at $T_{N}$, which we propose as the likely origin for the sharp peak in SHG that we observe. We emphasize that within this picture, the growth of SHG approaching $T_{N}$ results from an increase in the lifetime of magnetic fluctuations, not their amplitude. We can place a rough upper bound on the coherence time $\tau$ of the $4f^{7}\rightarrow 4f^{6}$ charge-transfer magnetic excitation. We observe an $O(1)$ change in SHG over the measured temperature range, which implies that $\gamma$ must be at least of the same order as $\Delta\omega$, the spectral width of the ultrafast laser. Thus, we calculate $\tau\sim 1/\gamma<1/\Delta\omega\sim 16$ fs. This value is comparable to the excitation lifetime of an isolated magnetic atom bound to a metallic substrate, as measured by inelastic tunneling spectroscopy, where it was suggested that the short lifetime was due to hybridization with metallic bands, leading to efficient electron-electron interactions that relax the magnetic state of the atom [34]. EuSn2(As,P)2 is also metallic, and it has been suggested that the magnetic exchange coupling in the material is mediated by conduction electrons via the Ruderman–Kittel–Kasuya–Yosida (RKKY) interaction [16, 35]. Indeed, the electrical conductivity of each compound directly correlates with $T_{N}$ (Supplementary Note 1), supporting a picture of carrier-mediated exchange. Thus, we may also expect a similarly short magnetic excitation lifetime due to coupling to conduction electrons.
@@
According to the fluctuation-dissipation theorem, there is a close relationship between the imaginary part of a linear susceptibility and the power spectral density of fluctuations of that susceptibility’s conjugate observable [33]. No such theorem exists for nonlinear susceptibilities, where the imaginary part does not represent dissipation [36]. It has been theoretically shown, however, that nonlinear responses are much more sensitive to semiclassical chaotic behavior through the interference of multiple dynamical Liouville-space trajectories [37]. The second-order susceptibility, for example, is directly related to the stability matrix of a system and its associated Lyapunov exponents quantifying the system’s sensitivity to initial conditions. Our measurements largely affirm this notion, demonstrating that, at least near resonances, the nonlinear optical response of a material can be significantly influenced by the dynamics of fluctuations of some order parameter of the material, even when that order is not the conjugate observable. We expect that this principle is generally applicable, and that under resonant conditions the nonlinear optical response can be harnessed to study fluctuations, dynamics, and quantum coherence across many types of phase transitions. In the case of magnetic systems, this information is complementary to that of the conventional linear optical susceptibility, whose real and imaginary parts have already been shown to be sensitive to lifetimes and entanglement of spin excitations in exotic magnets [38].
@@
Figure 3: Spontaneous rotational symmetry breaking in EuSn2As2. a Temperature dependence of the 2-fold rotational anisotropy that emerges at the antiferromagnetic phase transition. The inset shows Sin-Uout (SU) RA-SHG patterns for each temperature, and the main panel shows the corresponding 2-fold anisotropy versus temperature. The gray curve is a guide to the eye, showing an order-parameter-like onset below $T_{N}$. b Spatial map of rotational symmetry breaking. RA-SHG patterns were acquired on a $10\times 10$ grid spanning a $300\times 300$ $\mu$m area. At each grid point, the amplitude and direction of the 2-fold anisotropy was extracted (represented by the length and direction, respectively, of the double-headed arrows). The upper left inset shows a photograph of the sample with the mapped region identified, and the lower left inset shows a polar plot of the anisotropy sorted by direction into three symmetry-equivalent antiferromagnetic domains (red, green, and blue).
@@
Below $T_{N}$, the Eu2+ moments globally align, spontaneously breaking the $\bar{3}$ symmetry axis of the crystal as they choose to orient along one of three equivalent crystallographic directions. Rotational symmetry breaking can be revealed through RA-SHG patterns, which are exceptionally sensitive to point group symmetries [24]. As Fig. 3a demonstrates, we indeed observe rotational symmetry breaking below $T_{N}$, with a clear order-parameter-like onset. We measure the unpolarized SHG output (Uout) because a fully isotropic response is expected in the high-symmetry phase, degenerating into a 2-fold “peanut-like” shape when the symmetry axis is lost (Supplementary Note 5). We attribute this observation to two possible mechanisms, both involving spin-orbit coupling. In the first, coupling of the globally-aligned spins to the initial-state $f$ orbitals ($f_{xyz}$, $f_{xz^{2}}$, and $f_{x(x^{2}-3y^{2})}$) shifts their energies, and this manifests as slight changes to the optical resonance condition for different polarization angles, and in the second, global alignment of moments influences the quantum coherence of different initial-state $f$ orbitals. Accompanying spontaneous symmetry breaking is the formation of magnetic domains transforming into each other by the broken symmetry element. Optical imaging of antiferromagnetic domains is typically challenging because of the absence of a net magnetization, with SHG being one of the few known techniques for doing so [39, 40]. Figure 3b shows a spatial RA-SHG mapping over a $300\times 300$ $\mu$m area that reveals these domains. We find a roughly equal distribution of all three possible domain orientations, with a domain length scale of $\sim$100 $\mu$m. Thermal cycling across $T_{N}$ generates new domain patterns (Supplementary Note 6), suggesting domain boundaries are not strongly pinned by disorder.
III Discussion
@@
In conclusion, we harnessed near-resonant MD-SHG to probe quantum decoherence by magnetic fluctuations in the candidate axion insulator EuSn2(As,P)2. We observed an unusually sharp SHG peak at the magnetic phase boundary that we attribute to critical slowing down, where the dynamics of the system become extremely slow and the correlation length and time diverge. The quantum coherence time of the charge-transfer $4f^{7}\rightarrow 4f^{6}$ magnetic excitation is estimated to be no more than $\sim$16 fs. We conjecture that such a short lifetime is due to efficient coupling between the localized magnetic moments and the conduction electrons, supporting a scenario whereby RKKY interactions mediate magnetic exchange [16, 35]. This conclusion has important implications for the coupling strength between the magnetic order and the topological electronic states, where prior work has reported only a very weak hybridization [19]. We also uncovered evidence of strong short-range in-plane ferromagnetic correlations surviving far above $T_{N}$, which in similar materials has been proposed to stabilize topological phases [13] and has also been attributed to the possible formation of magnetic polarons [41]. Finally, we revealed the clear breaking of rotational symmetry within the antiferromagnetic phase and used it to map out the spatial domain structure. As a whole, our findings help to clarify the nature of the magnetic order in this class of materials and its interactions with topological surface states, which play a critical role in realizing topological magnetoelectric phenomena. More generally, our work shows that despite a lack of symmetry breaking enabling a direct coupling of the antiferromagnetic order parameter to SHG, resonant SHG can nevertheless still couple to magnetic fluctuations, providing access to antiferromagnetic order through an entirely different mechanism than considered by previous SHG studies of centrosymmetric antiferromagnets [42]. Ultimately, our work showcases the unique capabilities of nonlinear optical spectroscopy as a powerful tool for investigating quantum coherence, fluctuations, and symmetry breaking in magnetic quantum materials.
IV Methods
IV.1 Optical experiments
SHG experiments were performed on freshly-cleaved samples in an optical cryostat using a laser supplying 40 fs pulses at a 50 kHz repetition rate with a center wavelength of 800 nm and a fluence of $\sim$1 mJ/cm2. The oblique-incidence experimental geometry is illustrated in the upper left panel of Fig. 1c. The incidence angle relative to the sample surface normal was approximately 10. For the RA patterns shown in Fig. 1 and the temperature dependence shown in Fig. 2, the SHG signal was measured using a scientific CMOS camera. RA patterns were collected by directly rotating the sample, with the scattering plane remaining fixed. For the RA patterns and domain scans shown in Fig. 3, the SHG signal was measured using a silicon photomultiplier detector and a lock-in amplifier. For these measurements, normal incidence was used and the polarization of the incident beam was rotated instead of the sample, which afforded a much higher spatial resolution for mapping.
IV.2 Density functional theory calculations
DFT calculations were performed with the Vienna Ab initio Simulation Package (vasp) using the supplied projector augmented wave potentials within the generalized gradient approximation, the experimental structural parameters [16], a 400 eV energy cutoff, and a $6\times 6\times 6$ k-point mesh. Following prior work, to match the experimental energy position of the Eu $4f$ bands, a Hubbard parameter $U=5.0$ eV was employed [19].
IV.3 Sample synthesis and characterization
High quality single crystals of EuSn2As2 were prepared by a flux method similar to Ref. 21. Europium (99.9999%, Ames Lab) was cut up and combined with arsenic pieces (99.99%, Alfa Aesar) and Tin shot (99.9999%, Thermo Scientific) in a 1:3:20 molar ratio with a total mass of 4.0 g in an alumina crucible under a dry nitrogen glove box atmosphere. The crucible was sealed in a quartz ampoule under 300 Torr of argon without being exposed to air. The ampoule was heated in a box furnace from room temperature to 850C over 10 hours, held for 24 hours, then heated to 980C over 24 hours, before soaking at 980C for 12 hours. Finally, the ampoule was slowly cooled to 650C at 3.67C/hour, and spun in a centrifuge to remove excess tin flux. Shiny crystals in the form of hexagonal platelets with a typical diameter of 2 mm were obtained. Crystals of EuSn2P2 and EuSn2PAs were similarly prepared but with a different peak soaking temperature of 1000C and a centrifuging temperature of 700C, yielding crystals with a typical diameter of 1 mm. EuSn2As2 crystals grown without the heat soaks and the argon backfill pressure had notably smaller diameters, at around 200 $\mu$m. Both of these procedures have been used in previous growths of EuIn2As2 [22] to help avoid any high vapor pressure of arsenic and increase the homogeneity of the flux solution before reaching the peak temperature. We also note that growth attempts which centrifuged a Eu-Sn-As mixture at 750C yielded Eu5Sn2As6, a black needle-like crystal with a width of 100 $\mu$m and a length of a few mm. The structure and quality of the crystals was verified using powder x-ray diffraction on a Rigaku Ultima-4 system as well as energy-dispersive x-ray spectroscopy (EDX/EDS) on a SCIOS 2 DualBeam system.
V Data Availability
The data that support the findings of this study are available from the corresponding author upon reasonable request.
VI Acknowledgments
@@
VII Author Contributions
R.S., S.D., and R.S.R. conducted the optical measurements and analyzed the data. K.Y. and J.G.A. synthesized and characterized the samples. J.W.H. conceived and designed the project. R.S. and J.W.H. wrote the manuscript, with input from all authors.
VIII Competing Interests
The authors declare no competing interests.